OpenStudy (anonymous):

What is the value of the expression 3–4? A. –81 B. –12 C. -1/81 – D. 1/81

OpenStudy (anonymous):

3^(-4)=1/3^4=

OpenStudy (anonymous):

D!

OpenStudy (anonymous):

correct!